OpenStudy (anonymous):

How do you find the slope of the tangent line to the curve 2x^2+4xy+2y^3=114 at the point (-7,4) ??

OpenStudy (phi):

Do you know how to do "implicit differentiation" ? For a refresh, see https://www.khanacademy.org/math/differential-calculus/taking-derivatives/implicit_differentiation/v/implicit-differentiation-1 you will want to find dy/dx (as a function of x and y) using implicit differentiation. Evaluate at x=-7, y=4 to find the slope of the tangent line (i.e. the line you are looking for) If we call the slope m= dy/dx at (-7,4) find the equation of the line using \[ y - y_0 = m(x-x_0) \] where \( (x_0,y_0) = (-7,4) \)